OpenStudy (anonymous):

Given f(x) = the quantity of 3x minus 1, divided by 2 solve for f-1(4).

OpenStudy (gorv):

f(x)=(3x-1)/2 now to find inverse find value of x first

OpenStudy (gorv):

f=(3x-1)/2 2f=3x-1 2f+1=3x x=(2f+1)/3 this is ur inverse replce x by f^-1 and f by x f^-1=(2x+1)/3 now juxt put x=4 in it

OpenStudy (anonymous):

I'm getting -7/3 o_O, I don't think I calculated it properly. One second.

OpenStudy (anonymous):

Oh wow I put -4 instead of 4. Sorry

OpenStudy (anonymous):

So its 3?

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

Got 100% on the test. thank you Gorv!